OpenStudy (anonymous):

Manny and Emma are both competing in a long jump competition at their school. Manny can jump four fifths of the distance that Emma can jump. Manny can jump 40 inches. How far can Emma jump? A. 40.8 B. 50 C. 32 D. 39.2

OpenStudy (anonymous):

Manny = 4/5ths of Emma's height. Manny = 40 Emma = x Now create the equation using that info and solve for x.

OpenStudy (anonymous):

\[40={4\over5}x\]

OpenStudy (anonymous):

I was kinda hoping the creator of the question would do the equation themselves and learn the process.

OpenStudy (anonymous):

The equation is where the trouble was, I think. I could be wrong.

OpenStudy (anonymous):

I wasn't going to solve it :)

OpenStudy (anonymous):

I had problems setting it up,but thanks

OpenStudy (anonymous):

Don't forget your medal

OpenStudy (anonymous):

No one is required to give medals to anyone, the question asker doesn't have to. @dooley, when you have a question, I start like I did there, with what they gave you, the Manny = 4/5ths Emma's height. Then you put in what you know. They gave you manny's height of 40, so replace his name with that. Then they said to solve for emma, so you know that's your unknown, so you want to get it alone.

OpenStudy (anonymous):

No one is forcing lolz it was a friendly reminder, a lot of people forget the medal even exists.